OpenStudy (anonymous):

Simplify the expression: (-18 - 21r) ÷ (-12)

OpenStudy (ajprincess):

\[\frac{-18-21r}{-12}=\frac{-18}{-12}-\frac{21r}{-12}\] See if u can do it nw.

OpenStudy (jiteshmeghwal9):